Given:



1. Y=C+I+G+ X-M



2. C=b+cYd (Consumption function)



3. T=s+tY (tax function)



4. Mn + mY (import function)



5. Yd=Y-T (disposable income)



6. C = 100 + .90Yd



7. T = 40 + .20Y



8. M = 10 + .05Y



9. I = 38 G = 75 X = 25



Questions:



1. Calculate the equilibrium level of income using the model Y=C+I+G+X-M. 2. Determine the size of the new multiplier (import multiplier) for this economy.

1
Expert's answer
2021-12-14T15:05:33-0500

1.Y=C+I+G+X-M

Y=100 + 0.90Yd+38+75+25-(10 + 0.05Y)

Y=100 + 0.90(Y-T)+38+75+25-(10 + 0.05Y)

Y=100 + 0.90(Y-(40 +0.20Y))+38+75+25-(10 + 0.05Y)

Y=100+0.9(Y-40-0.2Y)+38+75+25-10-0.05Y

Y=100+0.9Y-36-0.18Y+38+75+25-10-0.05Y

Y=192+0.67Y

Y-0.67Y=192

0.33Y=192

Y=581.82

2.

"Me=\frac{1}{1-mpc+mpm}=\frac{1}{1-0.9+0.05}=6.67"